# taskpriors: Enhancing Model Evaluation by Considering the Entire Space of Downstream Tasks

### v2

@@ 3.3 SSL Objectives Can Serve as Task Priors
If you don’t have a prior kernel matrix, you may be able to use the same modeling assumptions you use in Self-Supervised Learning (SSL) to create an admissible kernel matrix. To start, while the above section demonstrates that supervised and SSL abide by the same objective, i.e., minimizing the absolute loss eq. 1 with ${\mathbf{Y}}$ or minimizing the relative loss eq. 2 with ${\mathbf{G}}={\mathbf{Y}}{\mathbf{Y}}^{\top}$ is equivalent, it remains to show what ${\mathbf{G}}$ SSL is actually employing.
@@
That is, SSL is implicitly acting on a labeling of the dataset that attributes to each sample a unique class. As such a labeling forces the model to maintain information about all its training samples–at least enough to separate them– it is clear that this is what maximized the Mutual Information between the original data ${\mathbf{X}}$ and the final learned representation $f_{{\mathbf{\theta}}}({\mathbf{X}})$ [20]. So, we can use this graph $\mathbf{G}$ as the kernel matrix to be our task prior.
3.4 Evaluating Model Kernels
We can use the equations in 3.6 as a way to evaluate model performance in a very fast way, i.e. without training any probes, or even assembling a collection of tasks / benchmarks. We demonstrate this in Figure 2 on a selection of models from timm and on a subset of $8,192$ images from mini-imagenet [19]. We use the centered cosine similarity as the choice of kernel matrix here and in the rest of the experiments in this paper. We find that the mean and variance are negatively correlated, implying that models that perform well on average tend also to perform better across a variety of tasks. From the selection of models we tested, we find that efficientnet [22] performs the best, even beating more modern vision transformers [6].
Figure 2: Here we plot the expectation and variance of $\operatorname{Tr}(GM)$, where $M$ is the centered cosine similarity kernel matrix for each models features generated from mini-imagenet[19], where the expectation is taken against $\mu_{K}$. Here we use $K$ as the centered cosine similarity kernel matrix for efficientnet_b5. Please see the appendix for more information and ablation on temperature and choice of kernel.
4 Applications of Task Priors
4.1 Empirically Sampling Tasks for Evaluation
@@
Sampling from the restricted measure $\mu_{K}^{q}$ is a much more challenging problem, and ends up being equivalent to sampling from the so-called Pott’s Model from statistical mechanics [28]. For instance, for binary labelings there are $2^{n}$ possible states, so sampling and computing the partition function can be completely intractable. We could use Markov Chain Monte Carlo (MCMC) methods such as the Metropolis-Hastings algorithm to sample from this distribution, but this can also prove to be challenging in practice and requires proper tuning of some parameters to get right. Instead, we propose an approximate sampling algorithm in $O(n)$ time to sample a labeling on $n$ data points.
@@
From (7), we can devise our method for the sampling Algorithm 1. Using this algorithm, we are able to quickly sample labelings of the data points according to the task prior, as demonstrated in Figure 1.
Key Takeaway: An $O(N)$ prefix-sampler lets us quickly realistic downstream tasks from the Task Prior, sidestepping costly MCMC and manual benchmark curation.
@@
1:$\mathbf{Z}\in\mathbb{R}^{n\times r}$triangleright factor so $\mathbf{K}\approx\mathbf{Z}\mathbf{Z}^{\top}$
2:$T>0$triangleright temperature
3:$q\in\mathbb{N}$triangleright number of classes
4:$\textit{labels}\in\{0,\dots,q-1\}^{n}$
@@
6:$\mathbf{U}\leftarrow 0_{r\times q}$triangleright class-wise prefix sums
@@
8:     $h\leftarrow(\frac{1}{T})\,(\mathbf{Z}[{i,:}]\;\mathbf{U})$triangleright length $q$ vector
9:     $h\leftarrow h-\max(h)$triangleright stabilize
10:     $p\leftarrow\exp(h)$;  $p\leftarrow p/\sum p$
@@
Figure 4: Correlation between the mean and variance of $\operatorname{Tr}(GK)$ and the performance of linear probes on the same representations, with $95\%$ confidence intervals.
4.2 Evaluating Models on Sampled Tasks
Using a selection of $33$ models from timm [27], we are able to evaluate each model on a distribution of tasks generated using efficientnet_b5 as our backbone prior. We report the results of this study in Figure 3, where we find that models that perform better on average also tend to have a better variance over tasks. This is a finding we will corroborate by directly measuring $\mathbb{E}_{\mu_{K}}[\operatorname{Tr}(GK)]$ and $\operatorname{Var}(\operatorname{Tr}(GK))$ in Figure 3.
@@ 4.2 Evaluating Models on Sampled Tasks
We note that the expectation and variance of $\operatorname{Tr}(GK)$ as shown in Figure 2, tends to exhibit the same trends as the models linear probe performance on sampled tasks, as seen in Figure 3, where stronger models tend to have a higher average accuracy / trace, as well as a lower variance. We compute the correlation between these sets of data points in Figure 4.
@@ 5 Discussion
Our results establish Task Priors as a principled bridge between intrinsic evaluations of representation learning models and the costly practice of curating ever-larger suites of downstream benchmarks. By casting the space of tasks as a distribution over label graphs, we obtain closed-form expressions for the expected performance and its variance across all possible tasks, and we show that these statistics align with empirical linear probe performance on ImageNet-derived data. Beyond evaluation, the theory suggests a prior-aware fine-tuning objective that simultaneously maximizes mean task performance while minimizing its dispersion.
@@ 5.1 Limitations and Future Work
Despite these advances, several open issues remain. While the trace metrics correlate with probe accuracy, the correlation is not exact; closing this theory–practice gap will require a deeper empirical and theoretic study. Additionally, storing the full $n^{2}$ kernel is can be prohibitive for very large datasets, although the matrices we observe are highly structured; further leveraging sparsity and low-rank factorizations is an immediate direction for further work. Our analysis is domain-agnostic, but its effectiveness on understanding the representations of Large Language Models [21], and, more generally, in natural language processing remains to be demonstrated. Our preliminary experiments hint that optimizing for the expected trace and its variance is beneficial, yet a full exploration of architectures, hyper-parameters, and interaction with existing self-supervised objectives is still outstanding. Tackling these questions will not only sharpen the foundations introduced here but may also lead to AI systems that perform consistently well across the vast landscape of tasks encountered in practice.
@@ Appendix A Main Theoretical Results
as desired. ∎
@@ Appendix B Additional Empirical Studies
Figure 5: Here, we show a comparison of how the choice of task prior kernel ${\mathbf{K}}$, reflected here in the color of the data points, affects the the evaluation of the mean and variance of $\operatorname{Tr}({\mathbf{M}}{\mathbf{G}})$. Each point is computed via the exact formulas given in 3.6, with a temperature of $T=0.01$.
@@ B.1 Ablation on Choice of Kernel
In Figure 5, we can see how the choice of task prior kernel matrix affects the downstream computation of the mean and variance of $\operatorname{Tr}({\mathbf{M}}{\mathbf{G}})$. As we might expect, we see that generally the mean $\mathbb{E}_{{\mathbf{G}}\sim\mu_{{\mathbf{K}}}}[\operatorname{Tr}({\mathbf{M}}{\mathbf{G}})]$ is higher when the task prior kernel ${\mathbf{K}}$ matrix is the same as the the matrix being evaluated ${\mathbf{M}}$. We don’t see this same behavior with the variance $\operatorname{Var}_{{\mathbf{G}}\sim\mu_{{\mathbf{K}}}}(\operatorname{Tr}({\mathbf{M}}{\mathbf{G}}))$.
@@ B.2 Ablation on the Temperature Parameter
Figure 6: Here we show a TSNE plot of Imagenette, with labels generated by the sampling Algorithm 1 for four choices of temperature.
Appendix C taskpriors PyTorch Package
Here we provide a brief demonstration of how one can use the provided codebase. We use the analyze function to compute the expectation and variance of a model with respect to a simple prior, as we compute in Theorem 3.6. We can also use the sampling method to sample labels as in Algorithm 1.
import torch
from torchvision import datasets, transforms, models
from torch.utils.data import DataLoader
from taskpriors import analyze
from taskpriors.sampler import sample_labels_from_model
# Select device
device = ’cuda’ if torch.cuda.is_available() else ’cpu’
# Data transformations: resize, crop, normalize
transform = transforms.Compose([
transforms.Resize(224),
transforms.CenterCrop(224),
transforms.ToTensor(),
transforms.Normalize([0.485, 0.456, 0.406],
[0.229, 0.224, 0.225]),
# Load Imagenette training split
dataset = datasets.Imagenette(root=’data’,
split=’train’,
transform=transform,
download=True)
loader = DataLoader(dataset,
batch_size=64,
shuffle=False)
# Initialize ResNet-18 backbone (remove final FC layer)
model = models.resnet18(
weights=models.ResNet18_Weights.IMAGENET1K_V1
model.fc = torch.nn.Identity()
model = model.to(device)
# Compute dataset statistics via TaskPriors
stats = analyze(model, loader)
print(stats)
# Sample soft labels from the model
labels = sample_labels_from_model(
model,
loader,
num_classes=10,
temperature=1000.0,
seed=0,
device=device
print(labels.shape)

### v3

@@ 2.3 SSL Objectives Can Serve as Task Priors
If you don’t have a prior kernel matrix, you may be able to use the same modeling assumptions you use in Self-Supervised Learning (SSL) to create an admissible kernel matrix (van den Oord et al., 2018; Bachman et al., 2019; He et al., 2020; Caron et al., 2020; Chen et al., 2020a; b; Henaff, 2020; Zbontar et al., 2021; Wang & Isola, 2020; Assran et al., 2023). To start, while the above section demonstrates that supervised and SSL abide by the same objective, i.e., minimizing the absolute loss eq. 1 with $\mathbf{Y}$} or minimizing the relative loss eq. 2 with ${\mathbf{G}}={\mathbf{Y}}{\mathbf{Y}}^{\top}$ is equivalent, it remains to show what $\mathbf{G}$} SSL is actually employing.
@@
That is, SSL is implicitly acting on a labeling of the dataset that attributes to each sample a unique class. As such a labeling forces the model to maintain information about all its training samples–at least enough to separate them– it is clear that this is what maximized the Mutual Information between the original data $\mathbf{X}$} and the final learned representation $f_{{\mathbf{\theta}}}({\mathbf{X}})$ (Shwartz-Ziv & LeCun, 2023). So, we can use this graph $\mathbf{G}$ as the kernel matrix to be our task prior.
Figure 3: We plot the expectation and variance of $\operatorname{Tr}(GM)$, where $M$ is the centered cosine similarity kernel matrix for each models features generated from mini-imagenet (Russakovsky et al., 2015), where the expectation is taken against $\mu_{K}$. Please see the appendix for more information and ablation on temperature and choice of prior kernel.
2.4 Sampling Tasks for Evaluation
@@
Sampling from the restricted measure $\mu_{K}^{q}$ is a much more challenging problem, and is equivalent to sampling from the Potts Model from statistical mechanics (Wu, 1982). For instance, for binary labelings there are $2^{n}$ possible states, so sampling and computing the partition function can be completely intractable. We could use Markov Chain Monte Carlo (MCMC) methods such as the Metropolis-Hastings algorithm to sample from this distribution, but this is sensitive to hyperparameters and may mix slowly in practice. Instead, we propose an approximate sampling algorithm in $O(n)$ time to sample a labeling on $n$ data points.
@@
From equation 7, we can devise our method for the sampling Algorithm 1. Using this algorithm, we are able to quickly sample labelings of the data points according to the task prior, as demonstrated in Figure 1, where we apply this on the Imagenette dataset (Howard, 2019) to features generated by the pretrained ResNet-18.  [^1] Interested readers can experiment with different class counts and temperature settings in this Colab notebook: https://colab.research.google.com/drive/1qNOgoNSH87AcdODug-yop7Q0MuT8w1r7
@@
[…]
2:$T>0$ $\triangleright$ temperature
[…]
@@
[…]
@@
8:  $h\leftarrow(\frac{1}{T})\,(\mathbf{Z}[{i,:}]\;\mathbf{U})$ $\triangleright$ length $q$ vector
9:  $h\leftarrow h-\max(h)$ $\triangleright$ stabilize
10:  $\leftarrow\exp(h);  pp$$\leftarrow p/\sum p$
@@
3 Empirical Applications of Task Priors
This section applies Task Priors to empirical evaluation of representation learning models. Here, we assess model kernels directly (3.1), show these statistics predict linear-probe accuracy on sampled labels (3.2), and validate alignment with curated benchmarks (3.3).
3.1 Evaluating Model Kernels
@@
Figure 4: Correlation between the mean and variance of $\operatorname{Tr}(GK)$, and the accuracy of linear probes sampled by the same Task Prior. We observe a Spearman correlation of $0.68$ (top) and $0.76$ (bottom).
We use the equations given in Theorem 2.6 as a way to evaluate model performance in a extremely efficient way, i.e. without training any probes, or even assembling a collection of tasks / benchmarks. We demonstrate this in Figure 3 and in Figure 2 (Left) on a selection of models trained to learn representations from images, on a subset of $8,192$ images from mini-imagenet (Russakovsky et al., 2015). We use the centered cosine similarity as the choice of kernel here and in the rest of the experiments in this paper. We also find that the mean and variance are negatively correlated, implying that models that perform well on average tend also to perform better across a variety of tasks.
3.2 Task Priors Predicts Performance on Sampled Labels
A central claim of this paper is that the two kernel statistics, $\mathbb{E}_{\mu_{K}}\!\left[\operatorname{Tr}(GK)\right]$ and $\operatorname{Var}_{\mu_{K}}\!\left(\operatorname{Tr}(GK)\right)$, can predict a representation’s downstream performance as measured by linear probes. For Figures 4 and 2 (Right), we sample tasks from a prior $\mu_{K}$ induced by DinoV2-giant. For Figure 4, we train an independent linear probe on every sampled task with and record the resulting accuracies. We then compare this to simply computing $\mathbb{E}_{\mu_{K}}[\operatorname{Tr}(GK)]$ and $\operatorname{Var}(\operatorname{Tr}(GK))$ as per Theorem 2.6.
@@ 3.2 Task Priors Predicts Performance on Sampled Labels
We note that the expectation and variance of $\operatorname{Tr}(GK)$ as shown in Figure 3, tends to exhibit the same trends as the model’s linear probe performance on sampled tasks, as corroborated by Figure 8, where stronger models tend to have a higher average accuracy / trace, as well as a lower variance. We observe a strong correlation between these sets of data points in Figure 4.
3.3 Task Priors Align with Performance on Curated Benchmarks
For this framework to be useful to practitioners, the performance on a hand-curated collection of downstream tasks should follow the distribution implied by the Task Prior. In the previous section section, we established that these theoretic measures based on the trace of the matrix ${\mathbf{G}}{\mathbf{K}}$ can represent performance on linear probes as generated via the task prior. In this section, we verify an analogous result but for a hand-curated collection of downstream tasks for image classification found in the MIEB (Xiao et al., 2025). We focus our analysis on a selection of $26$ models easily available through huggingface. The MIEB paper Tables $16$, $17$ report the accuracy of each of these models on each of $22$ downstream classification tasks. Using the strongest model on the MIEB tasks, google/siglip-so400m-patch14-384, as our task prior, we find that the mean and variance of the linear probe accuracy across these downstream tasks correlate to $\mathbb{E}_{\mu_{K}}[\operatorname{Tr}(GK)]$ and $\operatorname{Var}(\operatorname{Tr}(GK))$ respectively, as demonstrated in Figure 5. In this sense, Task Priors is able to predict the distribution of a model’s performance over real downstream tasks. We examine the bias incurred via the choice of prior model in Figure 6.
Figure 5: For each of the $26$ models, we estimate the task-prior expectation and variance and compare them to that model’s mean accuracy over a hand-curated set of $22$ MIEB classification tasks (Xiao et al., 2025). We observe a Spearman correlation of $0.82$ (Left) and $0.74$ (Right), showcasing how Task Priors can accurately predict model performance on a distribution of tasks.
4 Related Works
Many works aim to capture the performance of representation models primarily by intrinsic quantities about the model’s features. For instance, RankMe measures the effective rank of the feature matrix and shows a empirical correlation with average linear-probe accuracy across several tasks (Garrido et al., 2023). LiDAR argues that a method built on Linear Discriminant Analysis serves as a proxy for downstream performance (Thilak et al., 2023). More recent works take a broader view, demonstrating that k-NN, few-shot fine-tuning, and clustering evaluations may all disagree in systematic ways (Marks et al., 2025). Collectively, these studies show that properties intrinsic to the representation can forecast downstream success, but they still reduce performance to one or two scalar summaries.
A complementary line of work attacks the evaluation bottleneck by increasing the number of test tasks. In NLP, suites such as MTEB (56 embedding datasets) (Muennighoff et al., 2023) and HELM (42 scenarios, seven axes of measurement) (Liang et al., 2023) provide broad coverage of downstream tasks. The same trend is exists in vision, with works such as VideoEval packaging twenty diverse video understanding datasets together (Li et al., 2024), and frameworks such as MIEB (Xiao et al., 2025) providing a curated collection of downstream tasks for vision and multi-modal models.. While these large benchmarks can be quite helpful for practitioners, they remain finite and expensive to create. Worse, even a hundred benchmarks sample only a vanishingly small corner of the large task space practitioners can care about.
Our Task Priors framework can be viewed as the missing bridge between these two threads. Like intrinsic metrics, it avoids needing a hand curated set of downstream targets, but like conglomerate benchmarks, it explicitly reasons about many tasks. Section 2.1 follows closely the recent work of Balestriero & LeCun (2024). Our framework also echoes several well-known results from the classical theory of kernels. Notably, the trace term in Theorem 2.3 parallels the Hilbert–Schmidt Independence Criterion (HSIC) of Eq. (4) in (Gretton et al., 2007). Likewise, the term we get by taking the trace of ${\mathbf{K}}{\mathbf{G}}$ is precisely the same as “kernel alignment” studied in the context of generalization (Cristianini et al., 2001), obtained by flattening each matrix and taking their inner product, as $\langle{\mathbf{K}},{\mathbf{G}}\rangle=\operatorname{Tr}({\mathbf{K}}{\mathbf{G}})$.
There are some other related works that attack similar problems. In the computer vision space, VTAB (Zhai et al., 2020) takes a similar distributional view of tasks, but does not precisely characterize the distribution of tasks. Similar to our derivations, (Li et al., 2021) proposes a loss function based on the HSIC, which is an interesting avenue for future research.
@@ 5 Discussion
Our results establish Task Priors as a principled method to avoid the costly and time-consuming practice of curating ever-larger suites of downstream benchmarks. By casting the space of tasks as a distribution over label graphs, we obtain closed-form expressions for the expected performance and its variance across all possible tasks, and we show that these statistics align with empirical linear probe performance on data from ImageNet.
@@ 5.1 Limitations and Future Work
Despite these advances, several open issues remain. While the trace metrics correlate with probe accuracy, the correlation is not exact; closing this theory–practice gap will require a deeper empirical and theoretic study. Additionally, storing the full $n^{2}$ kernel can be prohibitive for very large datasets, although the matrices we observe are highly structured; further leveraging sparsity or sketching algorithms presents an immediate direction for further work. Our analysis is domain-agnostic, but our results in this paper focus on applications in learning representations for images. Further work is needed to examine its effectiveness on understanding the representations of Large Language Models (Skean et al., 2025), and, more generally, in natural language processing remains to be demonstrated. Additionally, this work focuses on a somewhat narrow definition of “task”, in that we focus on classification. Additional study should check if a variety of other tasks, e.g. retrieval, ranking, generative, segmentation, also agrees with the Task Prior, or if there are other ways to adopt a distribution over theses tasks. This work also suggests, a prior-aware fine-tuning objective that simultaneously maximizes mean task performance while minimizing its variance, which remains an open avenue for future research. Tackling these questions can help lead to AI systems that perform consistently well across the vast landscape of tasks encountered in practice.
5.1.1 Code and Reproducability
Code for this project can be found at: https://github.com/niketp03/taskpriors.
@@ Appendix A Main Theoretical Results
as desired. We can minimize over $\theta$ on the left hand side and right hand side of the above equation and we get our result. ∎
@@ B.1 Ablation on Choice of Kernel
In Figure 6, we can see how the choice of task prior kernel matrix affects the downstream computation of the mean and variance of $\operatorname{Tr}({\mathbf{M}}{\mathbf{G}})$. As we might expect, we see that generally the mean $\mathbb{E}_{{\mathbf{G}}\sim\mu_{{\mathbf{K}}}}[\operatorname{Tr}({\mathbf{M}}{\mathbf{G}})]$ is higher when the task prior kernel $\mathbf{K}$} matrix is the same as the the matrix being evaluated $\mathbf{M}$}. We can also see that models in the same family tend to demonstrate higher
Figure 6: Here, we show a comparison of how the choice of task prior kernel $\mathbf{K}$}, reflected here in the color of the data points, affects the the evaluation of the mean and variance of $\operatorname{Tr}({\mathbf{M}}{\mathbf{G}})$. Each point is computed via the exact formulas given in 2.6, with a temperature of $T=0.01$.
@@ B.2 Ablation on the Temperature Parameter
Figure 7: We show a TSNE plot of Imagenette, with labels generated by the sampling Algorithm 1 for four choices of temperature.
B.3 Evaluation of timm models with Task Priors
Using our sampling algorithm, we draw tasks from a prior $\mu_{K}$ induced by efficientnet_b5. For each of $33$ models from timm Wightman (2019), we train an independent linear probe on every sampled task with efficientnet_b5 and record the resulting accuracies. We then compare this to simply computing $\mathbb{E}_{\mu_{K}}[\operatorname{Tr}(GK)]$ and $\operatorname{Var}(\operatorname{Tr}(GK))$ as per 2.6.
We report the results of this study in Figure 8, where we find that models that perform better on average also tend to have a better variance over tasks.
Appendix C Code and Reproducibility
All code to reproduce the central results of this paper are available at the following anonymized repository: https://anonymous.4open.science/r/taskpriors-526A/